If a ring costs a jeweler $240, at what price should it be sold to yield a profit of 25% on the selling price?
Answer by stanbon(75887)

If a ring costs a jeweler $240, at what price should it be sold to yield a profit of 25% on the selling price?
----------
selling price = 240 + 0.25*240 = 1.25*240 = $300.00
===================================================
Cheers,
Stan H.
==========